OpenStudy (anonymous):

If f(x)=3x(2x+1)^(1/2) , what is f' (4) ?

OpenStudy (anonymous):

"f(x) = ... " is the classic way of writing a function.

OpenStudy (anonymous):

Yes, we would use the product rule here correct?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

1st take its 4th derivative and then put x=4 in 4th derivative equation , it is answer

OpenStudy (anonymous):

Derivative parts of f: \[f \prime = 3 (2x+1)^{1/2} + 3 x \frac{ 1 }{ 2 } (2x+1)^{-1/2}\] Then substitute x=4 into the above equation, you get the answer. \[f \prime (4) = 3.\sqrt{9} + 3. \frac{ 4 }{ 2. \sqrt{9} } = 11\]